# Primordial black hole capture, gravitational wave beats, and the nuclear equation of state

### v1

@@ Abstract
Primordial black holes (PBHs), if captured by neutron stars (NSs), would emit a characteristic gravitational wave (GW) signal as they orbit inside the host star. We identify a specific and qualitatively new feature of these signals, namely quasi-periodic beats caused by the precession of noncircular PBH orbits. We demonstrate numerically and analytically that the beat frequency depends rather sensitively on the NS structure, so that hypothetical future observations with next-generation GW detectors would provide valuable constraints on the nuclear equation of state.
@@
Primordial black holes (PBHs), first proposed by [1, 2], may have formed in the early Universe, and may contribute to or even make up most of its dark-matter content (see also [3]). While observational constraints on PBHs limit their possible contribution to the dark matter in some mass ranges, they remain viable candidates in other mass windows, including between about $10^{-16}M_{\odot}$ and $10^{-10}M_{\odot}$ as well as around $10^{-6}M_{\odot}$ (see, e.g., [4] for a review and details).
@@
If PBHs exist, some of them are likely to interact with stars and other celestial objects. Such interactions have been invoked as possible origins of several astrophysical phenomena, including the 1908 Tunguska event in Siberia [5] (but see [6]), neutron star (NS) implosions and “quiet supernovae” [7], fast radio bursts [8], the formation of low-mass stellar black holes [9], and the origin of supermassive stars (e.g. [10]). Gravitational-wave (GW) signatures of PBHs have been surveyed recently in [11], and the prospect of detecting PBHs using solar-system ephemerides has been discussed in [12, 13] and references therein.
@@
A collision with a star results in the PBH being gravitationally bound if it loses a sufficient amount of energy in the encounter, which is most likely to happen in collisions with NSs (see, e.g., [14, 8, 15, 16]). The PBH may still emerge from the star, but can no longer escape to infinity. Losing more energy in subsequent passages, the PBH at some point remains completely inside the star, settles down toward its center, accretes stellar material, and ultimately induces the dynamical collapse of the host star (see [17, 18, 19] for numerical simulations).
@@
While the PBH spirals toward the center of the NS it emits gravitational radiation that may be observable by next-generation GW detectors, and that would reveal information about the stellar structure. The authors of [20, 21], for example, examined this scenario assuming circular orbits. Since the PBH typically enters the host star on a noncircular orbit, and since the retarding forces inside the star may not circularize the orbit (see, e.g., [22, 23]), the PBH’s orbit is likely to be eccentric. In this letter we discuss a qualitatively new feature of such noncircular orbits, namely quasi-periodic gravitational-wave beats. These beats are caused by a precession of the PBH’s orbit inside the star, the GW frequency for which is superimposed on the higher frequency arising from a single orbit. The resulting GW envelopes for the two GW polarizations are exactly out of phase, so that the GW signal alternates between being dominated by one or the other polarization. In the stellar interior, both Newtonian and relativistic effects contribute to this precession, but we find that the latter dominate in NSs. As we demonstrate both numerically and analytically, the rate of the precession, and hence the beat frequency, depends rather strongly on the NS structure, so that a future observation of such a GW beat would provide strong constraints on the nuclear equation of state (EOS), let alone confirmation of the capture by a NS of a smaller and lower-mass intruder.
@@
As a numerical demonstration we show in Fig. 1 characteristic orbits of PBHs inside NSs governed by three different EOSs varying in stiffness, together with their associated GW signals. We adopt a simple particle test-mass approximation to describe the PBH moving on geodesics in the gravitational field of a relativistic star. As dynamical friction and accretion drag forces are small perturbations that operate on timescales much longer than orbital times, we can probe the precession by neglecting these forces and examining a few orbits via geodesics. We assume the stars to be governed by a polytropic EOS
@@
Here $P$ is the pressure, $\rho_{0}$ the rest-mass density, $K$ a constant, and the adiabatic exponent $\Gamma=1+1/n$ may be expressed in terms of the polytropic index $n$. We construct the stellar models by solving the Oppenheimer-Volkoff (OV) equations, adjusting the central density so that the stellar compaction is always given by $GM_{*}/(c^{2}R_{*})=1/6$, where $M_{*}$ is the star’s total gravitational mass and $R_{*}$ is areal radius (see Section I in the supplementary material for details). In Fig. 1 we show examples for a star with constant total mass-energy density $\rho$ (corresponding to $\Gamma=\infty$), as well as $\Gamma=3$ and $\Gamma=2$ polytropes, which serve as examples of both highly and moderately stiff candidates for the NS EOS.
Figure 1: Numerical examples of orbits inside NSs governed by different EOSs (left panels) together with the emitted GW signal as a function of time (right panels). The color shading in the left panels represents the density distribution inside the star. The top row shows, as the extreme limit, a constant-density star, corresponding to $\Gamma=\infty$. The middle and bottom rows show results for $\Gamma=3$ and $\Gamma=2$ polytropes. The GW amplitudes $h_{+}$ and $h_{\times}$ in the right panels are scaled for a star of mass $M_{*}=1.4M_{\odot}$ and hosting a black hole of mass $m=10^{-6}M_{*}$ at a distance of 10 kpc. The time $t$ in the right panel denotes time as measured by a distant observer, and is provided both in units of the the stellar mass $M_{*}$ (bottom axis) and in terms of ms (upper axis). An animated version of this figure is provided at [24].
We then solve the relativistic geodesic equations in order to track the PBH’s orbit. As discussed in Section II of the supplementary material, we choose to solve these equations in terms of the isotropic metric and radius $r$ rather than the interior Schwarzschild metric and areal radius $R$, although the conversion is straightforward. We always start orbits with vanishing radial speed $u^{r}=0$ at an initial (areal) radius $R(0)=R_{\rm frac}R_{*}$ and with angular momentum $\ell=\ell_{\rm frac}\ell_{\rm circ}$, where $\ell_{\rm circ}$ is the angular momentum corresponding to a circular orbit at radius $R(0)$. For the examples shown in Fig. 1 we used $R_{\rm frac}=0.6$ and $\ell_{\rm frac}=0.1$. All orbits are confined to a plane, which we arbitrarily take to be the $x-y$ plane. We also evaluate the leading-order GW signals $h_{+}$ and $h_{\times}$ along the $z-$axis using the quadrupole formalism.
@@
The above behavior can be understood in part in the context of Bertrand’s theorem, which states that closed orbits exist for potentials $V(R)=V_{0}+kR^{m}$, where $V_{0}$ and $k$ are constants, with exponents $m=-1$ and $m=2$ only.
The former exponent, $m=-1$, corresponds to a Newtonian point-mass potential, which is the leading order term in the potential in the exterior of a star. In general, deviations from this exterior potential result both from relativistic corrections as well as several Newtonian effects, including tidal and rotational deformations of the star (such deformations are not present for our static spherical stars). These deviations result in precession of the orbit, including the well-known relativistic perihelion advance of Mercury, which we reproduce in Section III.A of the supplemental material.
The latter exponent, a harmonic-oscillator potential with $m=2$, is realized in the interior of a Newtonian constant-density star, for which $M(R)=4\pi\rho R^{3}/3$ and hence $V(R)=V_{0}+2\pi G\rho R^{2}/3$. In this case, deviations result from both density nonuniformities and relativistic corrections. While we do not expect any precession, either in the interior or exterior, of a constant-density star in Newtonian gravitation, relativistic effects will cause precession for such a star both in the exterior and interior.
@@
In many cases (e.g., a Newtonian point-mass) $V(R)$ is independent of $e$ and $l$, but we now allow this term to depend on these two constants, which is the case for the orbits in general relativity considered here. In the following we assume the orbit to be in the equatorial plane, so that $\theta=\pi/2$, and we provide details in Section III of the Supplement of how $V(R)$ can be determined.
@@
While perturbations of the point-mass potential in the stellar exterior are familiar – yielding, for example, the relativistic perihelion advance of Mercury – we now focus on the stellar interior. Specifically, we show in Section III.B of the Supplement that, in the vicinity of the center, the potential $V(R)$ can be written in the form
@@
As the black hole advances from one periastron to the next, its positional angle $\varphi$ therefore advances by
@@
We refer the reader to Section III of the Supplement where the constants $p$ and $q$ are evaluated in general relativity for nearly circular orbits. Here we present a more transparent Newtonian treatment in order to illustrate the key ingredients. In the vicinity of the stellar center, we may approximate the density as $\rho(R)\simeq\rho_{c}+\rho^{(2)}R^{2}/2$, where $\rho^{(n})_{c}\equiv(d^{n}\rho/dR^{n})_{R=0}$. Integrating once we then find the enclosed mass $M(R)$, and integrating again we obtain the potential
@@
Comparing with (7) we can identify both $p$ and $q$ and compute
@@
Evidently, the Newtonian periastron advance is related to the degree of inhomogeneity, consistent with our discussion above.
@@
Table 1: Numerical and analytical data for polytropic stellar models and periastron advances for nearly circular orbits close to the center of a stellar host with compaction $GM_{*}/(c^{2}R_{*})=1/6$. We list, for different values of $\Gamma=1+1/n$, Newtonian values of the central condensation $\delta=\rho_{c}/\bar{\rho}$ (which depends on the $\Gamma$ alone) and the Newtonian estimate $\Delta\varphi/(R_{0}/R_{*})^{2}$, adopting numerical solutions to the Lane-Emden equation, together with $\Delta\varphi$ for $R_{0}/R_{*}=0.05$. For the relativistic data we computed $\delta$ from solutions to the OV equations. Adopting $R_{0}/R_{*}=0.05$ again we computed $\Delta\varphi_{\rm num}$ from numerical solutions to the geodesic equations, using $\ell_{\rm frac}=0.99$ for nearly circular orbits, and $\Delta\varphi_{\rm ana}$ analytically as presented in the Supplement. We also list the number of GW wave cycles ${\mathcal{N}}_{\rm GW}$ completed during a beat cycle (see Eq. 15), and, assuming a host star with mass $M_{*}=1.4M_{\odot}$, the orbital time $\tau_{\varphi}$ as well as the precession time $\tau_{\rm prec}$. In the last two columns we provide the corresponding GW frequencies $f_{\rm orb}^{\rm GW}\simeq 2(u^{t}\tau_{\varphi})^{-1}$ and $f_{\rm beat}^{\rm GW}\simeq(u^{t}\tau_{\rm prec})^{-1}$ as measured by a distant observer.
@@
where $a=(\Gamma P/\rho)^{1/2}$ is the (Newtonian) speed of sound. Finally we may use the central condensation $\delta\equiv\rho_{c}/\bar{\rho}$, where $\bar{\rho}=3M_{*}/(4\pi R_{*}^{3})$ is the average density, to rewrite the Newtonian periastron advance as
@@
For Newtonian polytropes $\delta$ depends on $\Gamma$ only (see Table 1 for specific values). For smaller $\Gamma$, $\delta$ is larger and $a_{c}^{2}$ smaller (for a given compaction $GM_{*}/(c^{2}R_{*})$); we therefore see that the periastron advance is larger for a softer EOSs (for a given value of $R_{0}/R_{*}$).
While the above analysis captures the leading-order Newtonian terms, we have found that the periastron advance in the NSs considered here is dominated by relativistic terms. However, the periastron advance’s dependence on the EOS’s stiffness is similar to that observed from the above Newtonian analysis even in the context of general relativity – namely, a softer EOS will lead to a more rapid precession of the orbit, and therefore to higher-frequency GW beats. This can be observed in Fig. 1 as well as in the Table 1, where we list periastron adnvances $\Delta\varphi$ for nearly-circular orbits ($\ell_{\rm frac}=0.99)$ close to the center ($R_{\rm frac}=0.05$) for a range of polytropic indices. We compare numerical results from the integration of the geodesic equation with analytical results from the perturbation of nearly-circular orbits and find excellent agreement.
We note that the range of polytropic exponents $\Gamma$ listed in Table 1 roughly covers values adopted in piecewise-polytropic approximations for candidate nuclear EOSs (see Table III in [25]; note in particular the larger range of values for $\Gamma_{3}$, which governs the high-density core). We find that the resulting values of $\Delta\varphi$ show significant variation, suggesting that a potential observation of the resulting GW beats would provide a sensitive probe of the EOS.
From the periastron advance $\Delta\phi$ we may also compute the precession frequency. Since we defined $\Delta\varphi$ as the (excess) advance from one periastron to the next, and since, to leading order, orbits in the stellar interior feature two periastrons per orbit, the angular precession frequency measured locally is given by $\Omega_{\rm prec}=2\Delta\varphi/\tau_{\varphi}$, where $\tau_{\varphi}$ is the orbital (proper) period. Related to the precession frequency is the (proper) precession period $\tau_{\rm prec}$ that it takes either of the two GW polarization amplitudes to go through one complete cycle, i.e. for the periastron to advance by an angle $\pi$, [^1] In the bottom panel of Fig. 1, for example, $\tau_{\rm prec}\simeq 1000M_{*}$.
@@
To summarize, we discuss GW beats as a qualitatively new feature of gravitational waves emitted by PBHs captured inside NSs. The beats are due to orbital precession, which is caused both by relativistic effects and density nonuniformity. Adopting a polytropic EOS we demonstrate both numerically and analytically that the beat frequency depends quite strongly on the structure of the NS and hence the stiffness of the EOS. For the NSs considered here the precession rate and beat frequency are largely due to relativistic gravitation, so that a Newtonian treatment would significantly underestimate the effect. If such beats were to be observed by next-generation GW detectors, e.g. the Einstein Telescope [26] or the Cosmic Explorer [27], they would therefore provide valuable constraints on the nuclear EOS.

### v2

# Primordial black holes, gravitational wave beats, and the nuclear equation of state

@@ Abstract
Primordial black holes (PBHs), if trapped in neutron stars (NSs), emit a characteristic continuous, quasiperiodic gravitational wave (GW) signal as they orbit inside the host star. We identify a specific and qualitatively new feature of these signals, namely quasiperiodic beats caused by the precession of noncircular PBH orbits. We demonstrate numerically and analytically that the beat frequency depends rather sensitively on the NS structure, so that hypothetical future observations with next-generation GW detectors could provide valuable constraints on the nuclear equation of state.
@@
Primordial black holes (PBHs), first proposed by [1, 2], may have formed in the early Universe, and may contribute to or even make up most of its dark-matter content (see also [3]). While observational constraints on PBHs limit their possible contribution to the dark matter in some mass ranges, they remain viable candidates in other mass windows, including between about $10^{-16}M_{\odot}$ and $10^{-10}M_{\odot}$ as well as around $10^{-6}M_{\odot}$ (see, e.g., [4, 5, 6] for reviews and details).
@@
If PBHs exist, some of them are likely to interact with stars and other celestial objects. Such interactions have been invoked as possible origins of several astrophysical phenomena, including the 1908 Tunguska event in Siberia [7] (but see [8]), neutron star (NS) implosions and “quiet supernovae” [9, 10], fast radio bursts [9, 11, 12], the formation of low-mass stellar black holes [13, 14, 15, 16], microquasars [17], and the origin of supermassive black holes (e.g. [18]), possibly via the formation of PBH clusters [19, 20]. Gravitational-wave (GW) signatures of PBHs have been surveyed recently in [21], and the prospect of detecting PBHs using solar-system ephemerides has been discussed in [22, 23] and references therein.
@@
A collision with a star results in the PBH being gravitationally bound if it loses a sufficient amount of energy in the encounter, which is most likely to happen in collisions with NSs (see, e.g., [24, 11, 25, 26]). The PBH may still emerge from the star, but can no longer escape to infinity. Losing more energy in subsequent passages, the PBH at some point remains completely inside the star, settles down toward its center, accretes stellar material, and ultimately induces the dynamical collapse of the host star (see [27, 28, 29] for numerical simulations). While the expected event rates are small (see, e.g., [30, 24, 25, 31, 32] as well as Section I in the supplementary material for estimates) they depend strongly on a number of assumptions and may be more favorable in special environments, e.g. globular clusters and galactic centers. Small black holes may also form inside neutron stars from the collapse of other dark-matter particles (e.g. [33, 34, 10, 27]), or be captured by neutron stars by other processes (e.g. [18, 35, 31, 26]).
@@
While the PBH spirals toward the center of the NS it emits gravitational radiation that – at least in principle – may be observable by next-generation GW detectors, and that would reveal information about the stellar structure [31]. The authors of [32, 36], for example, examined this scenario assuming circular orbits. Since the PBH typically enters the host star on a noncircular orbit, and since the retarding forces inside the star may not circularize the orbit (see, e.g., [37, 38]), the PBH’s orbit is likely to remain eccentric (see also [39] for a numerical demonstration). In this letter we discuss a qualitatively new feature of such noncircular orbits, namely continuous, quasiperiodic GW beats. These beats are caused by a precession of the PBH’s orbit inside the star, the GW frequency for which is superimposed on the higher frequency arising from a single orbit. The resulting GW envelopes for the two GW polarizations are exactly out of phase, so that the GW signal alternates between being dominated by one or the other polarization. In the stellar interior, both Newtonian and relativistic effects contribute to this precession, but we find that the latter dominate in NSs. As we demonstrate both numerically and analytically, the rate of the precession, and hence the beat frequency, depends rather strongly on the NS structure, so that a future observation of such a GW beat could provide strong constraints on the nuclear equation of state (EOS), let alone confirmation of the capture by a NS of a smaller and lower-mass intruder.
@@
As a numerical demonstration we show in Fig. 1 characteristic orbits of PBHs inside NSs governed by three different EOSs varying in stiffness, together with their associated GW signals. We adopt a simple particle test-mass approximation to describe the PBH moving on geodesics in the gravitational field of a relativistic star. As dynamical friction and accretion drag forces are small perturbations that operate on secular timescales much longer than orbital times (e.g. [30, 26]), we can probe the precession by neglecting these forces and examining a few orbits via geodesics. We assume the stars to be governed by a polytropic EOS
@@
Here $P$ is the pressure, $\rho_{0}$ the rest-mass density, $K$ a constant, and the adiabatic exponent $\Gamma=1+1/n$ may be expressed in terms of the polytropic index $n$. We construct the stellar models by solving the Oppenheimer-Volkoff (OV) equations, adjusting the central density so that the stellar compaction is always given by $GM_{*}/(c^{2}R_{*})=1/6$, where $M_{*}$ is the star’s total gravitational mass and $R_{*}$ is areal radius (see Section II in the supplementary material for details). In Fig. 1 we show examples for a star with constant total mass-energy density $\rho$ (corresponding to $\Gamma=\infty$), as well as $\Gamma=3$ and $\Gamma=2$ polytropes, which serve as examples of both highly and moderately stiff candidates for the NS EOS.
Figure 1: Numerical examples of orbits inside NSs governed by different EOSs (left panels) together with the emitted GW signal as a function of time (right panels). The color shading in the left panels represents the density distribution inside the star. The top row shows, as the extreme limit, a constant-density star, corresponding to $\Gamma=\infty$. The middle and bottom rows show results for $\Gamma=3$ and $\Gamma=2$ polytropes. The GW amplitudes $h_{+}$ and $h_{\times}$ in the right panels are scaled for a star of mass $M_{*}=1.4M_{\odot}$ and hosting a black hole of mass $m=10^{-6}M_{*}$ at a distance of $d=10$ kpc. The time $t$ in the right panel denotes time as measured by a distant observer, and is provided both in units of the the stellar mass $M_{*}$ (bottom axis) and in terms of ms (upper axis). (See [40] for an animation.)
We then solve the relativistic geodesic equations in order to track the PBH’s orbit. As discussed in Section III of the supplementary material, we choose to solve these equations in terms of the isotropic metric and radius $r$ rather than the interior Schwarzschild metric and areal radius $R$, although the conversion is straightforward. We always start orbits with vanishing radial speed $u^{r}=0$ at an initial (areal) radius $R(0)=R_{\rm frac}R_{*}$ and with angular momentum $\ell=\ell_{\rm frac}\ell_{\rm circ}$, where $\ell_{\rm circ}$ is the angular momentum corresponding to a circular orbit at radius $R(0)$. For the examples shown in Fig. 1 we used $R_{\rm frac}=0.6$ and $\ell_{\rm frac}=0.1$. All orbits are confined to a plane, which we arbitrarily take to be the $x-y$ plane. We also evaluate the leading-order GW signals $h_{+}$ and $h_{\times}$ along the $z-$axis using the quadrupole formalism [41, 42, 43].
@@
The above behavior can be understood in part in the context of Bertrand’s theorem ([44], see also [45]), which states that for potentials $V(R)=V_{0}+kR^{m}$, where $V_{0}$ and $k$ are constants, only the exponents $m=-1$ and $m=2$ will always result in closed orbits.
The former exponent, $m=-1$, corresponds to a Newtonian point-mass potential, which is the leading order term in the potential in the exterior of a star. In general, deviations from this exterior potential result both from relativistic corrections as well as several Newtonian effects, including tidal and rotational deformations of the star (which are not present for our static spherical stars). These deviations result in precession of the orbit, including the well-known relativistic perihelion advance of Mercury (see Section IV.A of the supplemental material).
The latter exponent, a harmonic-oscillator potential with $m=2$, is realized in the interior of a Newtonian constant-density star, for which $M(R)=4\pi\rho R^{3}/3$ and hence $V(R)=V_{0}+2\pi G\rho R^{2}/3$. In this case, deviations result from both density nonuniformities and relativistic corrections. While we do not expect any precession, either in the interior or exterior, for a homogeneous spherical star in Newtonian gravitation, relativistic effects will cause precession for such a star both in the exterior and interior.
@@
In many cases (e.g., a Newtonian point-mass) $V(R)$ is independent of $e$ and $l$, but we now allow this term to depend on these two constants, which is the case for the orbits in general relativity considered here. In the following we assume the orbit to be in the equatorial plane, so that $\theta=\pi/2$, and we provide details of how $V(R)$ can be determined in Section IV of the Supplement.
@@
While perturbations of the point-mass potential in the stellar exterior are familiar – yielding, for example, the relativistic perihelion advance of Mercury – we now focus on the stellar interior. Specifically, we show in Section IV.B of the Supplement that, in the vicinity of the center, the potential $V(R)$ can be written in the form
@@
As the black hole advances from one pericenter to the next, its positional angle $\varphi$ therefore advances by
@@
We refer the reader to Section IV of the Supplement where the constants $p$ and $q$ are evaluated in general relativity for nearly circular orbits. Here we present a more transparent Newtonian treatment in order to illustrate the key ingredients. In the vicinity of the stellar center we may approximate the density as $\rho(R)\simeq\rho_{c}+\rho^{(2)}R^{2}/2$, where $\rho^{(n})_{c}\equiv(d^{n}\rho/dR^{n})_{R=0}$. Integrating once we find the enclosed mass $M(R)$, and integrating again we obtain the potential
@@
Comparing with (7) we identify both $p$ and $q$ and compute
@@
Evidently, the Newtonian pericenter advance is related to the degree of inhomogeneity, consistent with our discussion above.
@@
Table 1: Numerical and analytical data for polytropic stellar models and pericenter advances for nearly circular orbits close to the center of a stellar host with compaction $GM_{*}/(c^{2}R_{*})=1/6$. We list, for different values of $\Gamma=1+1/n$, Newtonian values of the central condensation $\delta=\rho_{c}/\bar{\rho}$ (which depends on the $\Gamma$ alone) and the Newtonian estimate $\Delta\varphi/(R_{0}/R_{*})^{2}$, adopting numerical solutions to the Lane-Emden equation, together with $\Delta\varphi$ for $R_{0}/R_{*}=0.05$. For the relativistic data we computed $\delta$ from solutions to the OV equations. Adopting $R_{0}/R_{*}=0.05$ again we computed $\Delta\varphi_{\rm num}$ from numerical solutions to the geodesic equations, using $\ell_{\rm frac}=0.99$ for nearly circular orbits, and $\Delta\varphi_{\rm ana}$ analytically as presented in the Supplement. We also list the number of GW wave cycles ${\mathcal{N}}_{\rm GW}$ completed during a beat cycle (see Eq. 15), and, assuming a host star with mass $M_{*}=1.4M_{\odot}$, the orbital time $\tau_{\varphi}$ as well as the precession time $\tau_{\rm prec}$. In the last two columns we provide the corresponding GW frequencies $f_{\rm orb}^{\rm GW}\simeq 2(u^{t}\tau_{\varphi})^{-1}$ and $f_{\rm beat}^{\rm GW}\simeq(u^{t}\tau_{\rm prec})^{-1}$ as measured by a distant observer.
@@
where $a=(\Gamma P/\rho)^{1/2}$ is the (Newtonian) speed of sound. Finally we may use the central condensation $\delta\equiv\rho_{c}/\bar{\rho}$, where $\bar{\rho}=3M_{*}/(4\pi R_{*}^{3})$ is the average density, to rewrite the Newtonian pericenter advance as
@@
For Newtonian polytropes $\delta$ depends on $\Gamma$ only (see Table 1 for specific values). For smaller $\Gamma$, $\delta$ is larger and $a_{c}^{2}$ smaller (for a given compaction $GM_{*}/(c^{2}R_{*})$); we therefore see that the pericenter advance is larger for a softer EOSs (for a given value of $R_{0}/R_{*}$).
While the above analysis captures the leading-order Newtonian terms, we have found that the pericenter advance in the NSs considered here is dominated by relativistic terms. However, the pericenter advance’s dependence on the EOS’s stiffness is similar to that observed from the above Newtonian analysis even in the context of general relativity – namely, a softer EOS will lead to a more rapid precession of the orbit, and therefore to higher-frequency GW beats. This can be observed in Fig. 1 as well as in the Table 1, where we list pericenter adnvances $\Delta\varphi$ for nearly-circular orbits ($\ell_{\rm frac}=0.99)$ close to the center ($R_{\rm frac}=0.05$) for a range of polytropic indices. We compare numerical results from the integration of the geodesic equation with analytical results from the perturbation of nearly-circular orbits and find excellent agreement.
The range of polytropic exponents $\Gamma$ listed in Table 1 roughly covers values adopted in piecewise-polytropic approximations for candidate nuclear EOSs (see Table III in [46]; note in particular the larger range of values for $\Gamma_{3}$, which governs the high-density core). The resulting values of $\Delta\varphi$ show significant variation, suggesting that a potential observation of the resulting GW beats would provide a sensitive probe of the EOS.
From the pericenter advance $\Delta\phi$ we may also compute the precession frequency. Since we defined $\Delta\varphi$ as the (excess) advance from one pericenter to the next, and since, to leading order, orbits in the stellar interior feature two pericenters per orbit, the angular precession frequency measured locally is given by $\Omega_{\rm prec}=2\Delta\varphi/\tau_{\varphi}$, where $\tau_{\varphi}$ is the orbital (proper) period. Related to the precession frequency is the (proper) precession period $\tau_{\rm prec}$ that it takes either of the two GW polarization amplitudes to go through one complete cycle, i.e. for the pericenter to advance by an angle $\pi$, [^1] In the bottom panel of Fig. 1, for example, $\tau_{\rm prec}\simeq 1000M_{*}$.
@@
To summarize, we discuss quasiperiodic GW beats as a qualitatively new feature of continuous GW signals emitted by PBHs captured inside NSs. The beats are due to orbital precession, which is caused both by relativistic effects and density nonuniformity. Adopting a polytropic EOS we demonstrate both numerically and analytically that the beat frequency depends quite strongly on the structure of the NS and hence the stiffness of the EOS. For the NSs considered here the precession rate and beat frequency are largely due to relativistic gravitation, so that a Newtonian treatment would significantly underestimate the effect. If such beats were to be observed by next-generation GW detectors, e.g. the Einstein Telescope [47], the Cosmic Explorer [48], or the Neutron Star Extreme Matter Observatory (NEMO) [49], they would therefore provide valuable constraints on the nuclear EOS.